bill533:

What goes in the question ? file:///C:/Users/user/Pictures/s5FJ7uT.jpg

Gdeinward:

Hey man, you gotta find another way to post. My computer doesnt allow me to go to imgur

bill533:

alright hold on

bill533:

how about now

Gdeinward:

same thing. Can you take a screen shot?

bill533:

i did on these

Gdeinward:

I guess you will have to post the old fashioned way, sorry man.

bill533:

i guess so bruh

Gdeinward:

I could acess it form my phone.

bill533:

sure

bill533:

i just sent it

Gdeinward:

its 3^4 * 5 * 7

Gdeinward:

heres a link that explains it: https://www.2dtx.com/prime/prime2835.html

bill533:

i have a couple more if you would

Gdeinward:

sure

bill533:

i just sent it

Gdeinward:

im solving it givre me a sec

bill533:

yes sir

Gdeinward:

hey can you give me the numbersin the first matrix?

Gdeinward:

starting from left to right going down

Gdeinward:

so I have 2 0 1 ? 1 9

Gdeinward:

I cant read what the number is in the place of the question mark

bill533:

2 0 8 -3 1 9

Gdeinward:

thank you

bill533:

no prob

Gdeinward:

sorry dude, the answer is C. I solved it, heres a link to a claculater that describes how to solve matrix operations.

Gdeinward:

https://matrixcalc.org/

Gdeinward:

it is not good to use a calulater, because you will never learn it, so try to do these problems on your own.

Gdeinward:

Then check them with the calculater to learn from your mistakes

bill533:

ok another one

Gdeinward:

its B. he created the equation to determine the height after two weeks.

Gdeinward:

D is not defined

bill533:

i left another one but you know what look at your gmail everytime i sent you something

Gdeinward:

its D 9/29. 45/ 5 = 9 145/5 = 29

Gdeinward:

E.

Gdeinward:

just post it on here. I cant help with anymore sorry. I have Algebra I need to work on.

Gdeinward:

I will notify Dude so you can get the help you need.

bill533:

it's 2 last question after this and it's over

Gdeinward:

or somebody else

bill533:

i appreciate ur help man but if u help with 2 questions after i'll be done

Gdeinward:

Its D. here is a link to a website calculater that shows step by step how to solve. https://www.symbolab.com/solver/step-by-step/6%20-%202i%20%2F%205%20%2B%203i

bill533:

i appreciate a lot for your help man, after these questions that's it

Gdeinward:

I beleive the first on the answer is A.

Gdeinward:

Im stumped on that last one.

Gdeinward:

Its the only one I dont know how to solve

bill533:

for the first one, is that right

Gdeinward:

Im confused. the calculate works so that is the solution. I dont recall you sending me that problem

Gdeinward:

on that last one you sent me I found out that the product of matrices is ASSOCIATIVE, that is for matrices A, B, C, we have: A*B*C=(A*B)*C=A*(B*C)

Gdeinward:

I cant help anymore than that. sorry.

bill533:

ok thanks a lot

Gdeinward:

your welcome.